(D ^ 2 + 3D + 2) * y = sin 2x + 5 degrees + log 3

1
Expert's answer
2022-01-31T16:38:05-0500

The given question is;

"\displaystyle
(D^2+3D+2)y=\sin(2x+5)+\log3"

The Auxiliary equation of the given DE in terms of variable m is;

"\displaystyle
m^2+3m+2=0\\"

"\displaystyle
\Rightarrow(m+1)(m+2)=0\\
\Rightarrow m=-1,-2\\"

Hence, the complementary function is;

"\displaystyle
y_c=c_1e^{-x}+c_2e^{-2x}", where "\displaystyle
c_1" and "\displaystyle
c_2" are arbitrary constants.


Next, to calculate the particular integral "\displaystyle
(y_p)" we proceed as follows;

"\displaystyle
y_p=\frac{1}{D^2+3D+2}\sin2\left(x+\frac{5}{2}\right)+\frac{1}{D^2+3D+2}(\log3)\\
\quad=\frac{1}{-4+3D+2}\sin2\left(x+\frac{5}{2}\right)+\log3\frac{1}{D^2+3D+2}e^{0x}\\
\quad=\frac{1}{3D-2}\sin2\left(x+\frac{5}{2}\right)+\frac{\log3}{2}\\
\quad=\frac{3D+2}{9D^2-4}\sin2\left(x+\frac{5}{2}\right)+\frac{\log3}{2}\\
\quad=\frac{3D+2}{-40}\sin2\left(x+\frac{5}{2}\right)+\frac{\log3}{2}\\
\quad=\frac{-1}{40}\left[3D\left(\sin(2x+5)\right)+2\sin(2x+5)\right]+\frac{\log3}{2}\\
\quad=\frac{-1}{40}[6\cos(2x+5)+2\sin(2x+5)]+\frac{\log3}{2}\\
\quad=-\frac{3\cos(2x+5)}{20}-\frac{\sin(2x+5)}{20}+\frac{\log3}{2}"


Hence, the general solution is;

"\displaystyle
y=y_c+y_p\\
\quad=c_1e^{-x}+c_2e^{-2x}-\frac{3\cos(2x+5)}{20}-\frac{\sin(2x+5)}{20}+\frac{\log3}{2}"